Joseph Mart recently purchased new trucks for its delivery fleet. The truck manufacturer confirmed

that the engines in the trucks have a mean failure time of 8 years, with a standard deviation of

1.5 years. Within what amount of time should Walmart expect for 93.32% of the trucks to fail?

Please show all work, with a normal distribution curve if possible, thanks in advance!
Answer #1

Solution:

Given: the trucks have a mean failure time of 8 years, with a standard deviation of 1.5 years.

That is : Mean = u= 8 and standard deviation = 0 = 1.5

We have to find the values of x such that:

P( x1 < X < x2 ) = 93.32%

P( x1 < X < x2 ) = 0.9332

Thus find z values such that:

P( -z  < Z < z  ) = 0.9332

If area between -z to +z is 0.9332 , then area outside this interval is = 1 - 0.9332 = 0.0668

That is area in tails is 0.0668

We divide this area in two tails equally

that is: 0.0668 / 2 = 0.0334 in left and right tail.

That is:

P( Z< -z) = 0.0334  

and

P( Z> z )= 0.0334

Thus we get:

Pl-z<Z<z )= 0.9332 P(Z > z) = 0.0334 P(Z <-z) = 0.0334

Thus from above graph we can see total area below z is:

P( Z< z ) = P( Z< -z ) + P( -z < Z < z )

P( Z< z ) = 0.0334 + 0.9332

P( Z< z ) = 0.9666

Thus look in z table for area = 0.9666 or its closest area and find z value.

10.6 1.00 .01 .02 .4 .05 .06 .07 .08 .09] | 0.0 - 1,5000 5040 5080 .50 5160 .5199 .5239 5279 5319 .5359 | 10.1 ,5398 ,5438 .5

Area 0.9664 is closest to 0.9666 and it corresponds to 1.8 and 0.03

thus z = 1.83

thus -z = -1.83

Now for finding x value we use formula:

= μ +Χσ

T1 = 1+(-x) X Ở

I1 = 8+(-1.83) x 1.5

-21 = 8 -2.745

21 = 5.255

-21 = 5.3

and

T2 = μ+ (+2) xσ

22 = 8+1.83 x 1.5

x_{2} = 8 +2.745

C2 = 10.745

12 = 10.7

Thus within 5.3 years to 10.7 years   should Walmart expect for 93.32% of the trucks to fail.

P(5.3 < X < 10.7) = 0.9332 -2 -1.83 x1 = 5.3 years mean=8 years z = 1.83 x2 = 10.7 years